1- On July 1, 2018, Tremen Corporation acquired 35% of the shares of Delany Company. Tremen paid $3,060,000 for the investment, and that amount is exactly equal to 35% of the book value of identifiable net assets on Delany's balance sheet. Delany recognized net income of $1,000,000 for 2018, and paid $180,000 of dividends each quarter to its shareholders. After all closing entries are made, Tremen's "Investment in Delany Company" account would have a balance of:

2-On January 1, 2018, Green Corporation purchased 34% of the outstanding voting common stock of Gold Company for $300,650. The book value of the acquired shares was $275,650. The excess of cost over book value is attributable to an intangible asset on Gold's books that was undervalued and had a remaining useful life of five years. For the year ended December 31, 2018, Gold reported net income of $125,550 and paid cash dividends of $25,900. What is the carrying value of Green's investment in Gold at December 31, 2018?

3-On December 31, 2018, L Inc. had a $2,000,000 note payable outstanding, due July 31, 2019. L borrowed the money to finance construction of a new plant. L planned to refinance the note by issuing long-term bonds. Because L temporarily had excess cash, it prepaid $550,000 of the note on January 23, 2019. In February 2019, L completed a $3,500,000 bond offering. L will use the bond offering proceeds to repay the note payable at its maturity and to pay construction costs during 2019. On March 13, 2019, L issued its 2018 financial statements. What amount of the note payable should L include in the current liabilities section of its December 31, 2018, balance sheet?
Answer #1

1.

The period is July 1 to Dec 31 = 6 months

The balance of investment = 3060,000 + 35% of (1000,000 – 180,000 X 4) X 6/12months

                                                                = 3060,000 + 49,000 = $ 3,109,000

2.

Cost

$        300,650

Share of Net income (125,550 X 34%)

$          42,687

Share of dividends (25,900 X 34%)

$          (8,806)

Amortization of intangibles (300,650- 275,650)/5

$          (5,000)

Carrying value

$        329,531

Hence, carrying value of Green's investment in Gold at December 31, 2018 is $ 329,531.

3.

As per GAAP, the amount excluded from current liabilities through refinancing cannot exceed the actual amount refinanced.

In the given case, L should opt for paid 2000,000 – 550,000 =$ 1,450,000 to be long term liability and the balance of $ 550,000 as current liability. As the refinance arrangement is done before issuance of financial statements, there is both intent and ability on part of L.

Hence, L should include $ 550,000 as part of notes payable as current liabilities

kindly upvote